The company with the common equity accounts shown here has declared a 4-for-one stock split when the market value of its stock is $35 per share. The firm’s 60-cent per share cash dividend on the new (postsplit) shares represents an increase of 25 percent over last year’s dividend on the presplit stock.

  

  Common stock ($1 par value) $ 485,000
  Capital surplus 866,000
  Retained earnings 3,920,800
     Total owner's equity $ 5,271,800

What is the new par value per share? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.)

  New par value $  per share  

What was last year's dividend per share? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.)

  

  Dividend per share $   
Answer #1

1. New par value = Current par value / No of shares in exchange = 1/4 = $0.25

2. Dividend per share = ( % of cash dividend * 4) / ( 1+25%) = ( 0.6* 4) / (1.25) = $1.92 per share